# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.1. The affine Hecke algebra and simple supersingular ${\mathcal{H}}$-modules
It remains to show that ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{\chi}}({\mathcal{H}})$ is (left and right) Noetherian. Write $\Omega_{\chi}$ for the image of $\widetilde{\Omega}_{\chi}$ in $\Omega$. This group decomposes as $\Omega_{\chi}=\Omega_{\chi}^{\text{free}}\times\Omega_{\chi}^{\text{tor}}$ where $\Omega_{\chi}^{\text{free}}$ is a free abelian group on some generators $\omega_{1},\ldots,\omega_{r_{Z}}$ and $\Omega_{\chi}^{\text{tor}}$ is finite. If we let ${\mathcal{H}}_{\chi}^{\text{free}}$ denote the subalgebra of ${\mathcal{H}}_{\chi}$ generated by ${\mathcal{H}_{\operatorname{aff}}}$ and all $T_{\hat{\omega}_{i}^{\pm 1}}$, then the braid relations show that ${\mathcal{H}}_{\chi}^{\text{free}}$ is an iterated skew Laurent polynomial algebra over ${\mathcal{H}_{\operatorname{aff}}}$, and hence is Noetherian because ${\mathcal{H}_{\operatorname{aff}}}$ is Noetherian (c.f. [32, Theorem 5.1] and [21, Theorem 1.4.5]). Since ${\mathcal{H}}_{\chi}$ is finite over ${\mathcal{H}}_{\chi}^{\text{free}}$, it must be Noetherian as well. ∎
We now suppose that $\chi=(J,\xi)$ is a supersingular character of ${\mathcal{H}_{\operatorname{aff}}}$. Given a finite dimensional, irreducible (right) representation $(V,\rho)$ of $\widetilde{\Omega}_{\chi}$ on which $T(\mathbb{F}_{q})$ acts by the character $\xi$, we let $\chi\otimes V$ be the ${\mathcal{H}}_{\chi}$-module whose underlying vector space is $V$ and on which ${\mathcal{H}_{\operatorname{aff}}}$ acts by $\chi$ and $\widetilde{\Omega}_{\chi}$ acts via $\rho$. Then every simple supersingular ${\mathcal{H}}$-module is isomorphic to one of the form $(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$, and furthermore two pairs $(\chi,V)$ and $(\chi^{\prime},V^{\prime})$ give rise to isomorphic simple supersingular ${\mathcal{H}}$-modules if and only if the two pairs are $\widetilde{\Omega}$-conjugate (c.f. [32, Theorem 6.18]). Here we say that $(\chi,V)$ and $(\chi^{\prime},V^{\prime})$ are $\widetilde{\Omega}$-conjugate if $\chi^{\prime}=\chi^{\tilde{\omega}}$ for some $\tilde{\omega}\in\widetilde{\Omega}$ and $V^{\prime}\cong V^{\tilde{\omega}}$, where $V^{\tilde{\omega}}$ is the representation of $\widetilde{\Omega}_{\chi^{\tilde{\omega}}}=\tilde{\omega}^{-1}\widetilde{\Omega}_{\chi}\tilde{\omega}$ with underlying vector space $V$ obtained by letting $v\cdot\tilde{\omega}^{-1}\tilde{\omega}^{\prime}\tilde{\omega}:=v\cdot\tilde{\omega}^{\prime}$ for all $v\in V$ and $\tilde{\omega}^{\prime}\in\widetilde{\Omega}_{\chi}$.
@@
We finally state a theorem of Koziol, which describes when simple supersingular modules are trivial in the Gorenstein projective model structure. The statement in loc. cit. doesn’t actually contain (ii) as below, but their proof actually shows the statement as written here.
Theorem 5.4 ([19, Theorem 7.7]).
@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
Our first aim is to understand morphisms between supersingular characters in the homotopy category ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. Fix a face $F\subseteq\overline{C}$. Recall that ${\mathcal{H}}_{F}$ is a Frobenius algebra and in particular selfinjective (c.f. [9, Example 4.1(ii)]), and thus the Gorenstein projective model structure on ${\operatorname{Mod}}({\mathcal{H}}_{F})$ has all objects cofibrant and fibrant. In particular, we see from 2.14 that
@@
where two morphisms $f$, $g$ are equivalent under $\sim$ if and only if $f-g$ factors through a projective ${\mathcal{H}}_{F}$-module. We collect the various facts we need about the interactions between the Gorenstein projective model structures on the categories ${\operatorname{Mod}}({\mathcal{H}}_{F})$, $F\subseteq\overline{C}$, and ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ below:
@@
[…]
@@
    is right Quillen. Moreover, the Gorenstein projective model structure on ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ is the right transfer of the product model structure on $\prod_{F\subseteq\overline{C}}{\operatorname{Mod}}({\mathcal{H}}_{F})$ (i.e. with all classes of morphisms defined termwise) along $\Delta$.
@@
For (i), we have that $[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}=0={\operatorname{Hom}}_{{\mathcal{H}}_{F}}({\mathfrak{m}},\mathfrak{n})$ if the simple modules ${\mathfrak{m}}$ and $\mathfrak{n}$ are not isomorphic (c.f. (5.4)). If on the other hand ${\mathfrak{m}}\cong\mathfrak{n}$, then by Schur’s lemma we have that ${\operatorname{Hom}}_{{\mathcal{H}}_{F}}({\mathfrak{m}},{\mathfrak{m}})\cong k$ is one-dimensional and hence either its quotient $[{\mathfrak{m}},{\mathfrak{m}}]_{{\mathcal{H}}_{F}}$ is zero or the homotopy relation is trivial. But if $[{\mathfrak{m}},{\mathfrak{m}}]_{{\mathcal{H}}_{F}}=0$ then ${\operatorname{id}}:{\mathfrak{m}}\to{\mathfrak{m}}$ factors through a projective, from which we deduce that ${\mathfrak{m}}$ is projective. We thus get our desired isomorphism.
Part (ii) is immediate from 2.13. For (iii), we obtain as a consequence of (ii) that $\Delta$ preserves fibrations and weak equivalences as those are defined termwise in the target category. Since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}$ is also faithful for each $F\subseteq\overline{C}$, we in fact get that $\Delta$ reflects fibrations. We are therefore left to show that $\Delta$ has a left adjoint and reflects weak equivalences. For the former, the left adjoint is given by $({\mathfrak{m}}_{F})_{F\subseteq\overline{C}}\mapsto\bigoplus_{F\subseteq\overline{C}}{\mathfrak{m}}_{F}\otimes_{{\mathcal{H}}_{F}}{\mathcal{H}_{\operatorname{aff}}}$. For the latter, by 2.8 it suffices to show that $\Delta$ reflects trivial objects, i.e. that if ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ is such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$ is projective for all $F\subseteq\overline{C}$ then ${\operatorname{pd}}_{{\mathcal{H}_{\operatorname{aff}}}}({\mathfrak{m}})<\infty$. But this is the reverse implication shown in [19, Lemma 4.2]. ∎
5.5(i) has the following immediate consequence which will be helpful later.
@@
    $(\chi^{\prime}(T_{\hat{s}}),\chi^{\prime}(T_{\hat{s^{\prime}}}),\chi^{\prime}(T_{\hat{s^{\prime\prime}}}))$ $=(0,-1,0).$
    Then ${\operatorname{Ho}}(\Delta)([\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}})$ has dimension 1 over $k$ and doesn’t contain any isomorphisms.
@@
Later we will consider simple supersingular modules for the whole pro-$p$ Iwahori-Hecke algebra ${\mathcal{H}}$, for which the action of $\widetilde{\Omega}$ on supersingular characters comes into play. Suppose that $\omega\in\Omega$ and fix a lift $\hat{\omega}\in\widetilde{\Omega}$. Recalling that the group $\Omega$ normalises $S$, if $\chi$ is any character of ${\mathcal{H}_{\operatorname{aff}}}$ corresponding to a pair $(J,\xi)$ then $\chi^{\hat{\omega}}$ corresponds to the pair $(J^{\hat{\omega}},\xi^{\hat{\omega}})$ where $J^{\hat{\omega}}=\omega J\omega^{-1}$ and $\xi^{\hat{\omega}}(t)=\xi(\hat{\omega}t\hat{\omega}^{-1})$ ($t\in T(\mathbb{F}_{q})$). It follows in particular that $\left|J^{\hat{\omega}}\right|=\left|J\right|$. We therefore see that the two supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ from 5.7(i) are not conjugate under the $\widetilde{\Omega}$-action.
We start preparing for the proof. Suppose that $\chi$ and $\chi^{\prime}$ are as given in 5.7. Given any $f\in[\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}}$ we obtain a collection $(f_{F})\in\prod_{F\subseteq\overline{C}}{\operatorname{Hom}}_{{\mathcal{H}}_{F}}(\chi,\chi^{\prime})$ by letting each $f_{F}$ be a choice of representative of ${\operatorname{Ho}}(\Delta)(f)_{F}={\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}})(f)$, and this satisfies the property that ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F}}(f_{F^{\prime}})\sim f_{F}$ whenever $F^{\prime}\subseteq\overline{F}$. Note that either $f_{F}\sim 0$ or otherwise the choice of $f_{F}$ must be unique, c.f. 5.5(i). The strategy for our proof will be to show more generally that any such collection of morphisms must in fact be zero, up to homotopy, unless we are in the specific circumstances of part (i) of the Proposition. We therefore first need a method for determining whether some $f_{F}$ arising as above is homotopic to zero. To that end, we have the following elementary but key result.
@@
    If we have three faces $F_{1},F_{2},F_{3}$ with $F_{i}\subseteq\overline{F_{3}}$ for $i=1,2$ such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective and $f_{F_{2}}\sim 0$ then $f_{F_{1}}\sim 0$.
@@
Part (i) is immediate from 5.5(ii) because right Quillen functors preserve right homotopies. For part (ii), assume that $f_{F}\sim 0$ and $f_{F^{\prime}}\not\sim 0$. Then by 5.5(i) we must have ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ and $f_{F^{\prime}}=\lambda{\operatorname{id}}_{\chi}$ for some $\lambda\in k^{\times}$.Then we have ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\lambda{\operatorname{id}}_{\chi})\sim f_{F}\sim 0$ from which we deduce that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective. For (iii), the condition that $f_{F_{2}}\sim 0$ implies by (i) that $f_{F_{3}}\sim 0$. Since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective it then follows from (ii) that $f_{F_{1}}\sim 0$ as required. ∎
In order to apply 5.9, we need some criterion to establish when the restriction of a character of ${\mathcal{H}_{\operatorname{aff}}}$ to some ${\mathcal{H}}_{F}$ is not projective. For this, we have the result below which we gathered from the arguments of Koziol in [19, §7].
@@
    If $S_{F}\cap S_{\xi}\neq S_{F}$ then ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective.

### v4

@@ 5.1. The affine Hecke algebra and simple supersingular ${\mathcal{H}}$-modules
It remains to show that ${\mathcal{H}}_{\chi}$ is (left and right) Noetherian. Write $\Omega_{\chi}$ for the image of $\widetilde{\Omega}_{\chi}$ in $\Omega$. This group decomposes as $\Omega_{\chi}=\Omega_{\chi}^{\text{free}}\times\Omega_{\chi}^{\text{tor}}$ where $\Omega_{\chi}^{\text{free}}$ is a free abelian group on some generators $\omega_{1},\ldots,\omega_{r_{Z}}$ and $\Omega_{\chi}^{\text{tor}}$ is finite. If we let ${\mathcal{H}}_{\chi}^{\text{free}}$ denote the subalgebra of ${\mathcal{H}}_{\chi}$ generated by ${\mathcal{H}_{\operatorname{aff}}}$ and all $T_{\hat{\omega}_{i}^{\pm 1}}$, then the braid relations show that ${\mathcal{H}}_{\chi}^{\text{free}}$ is an iterated skew Laurent polynomial algebra over ${\mathcal{H}_{\operatorname{aff}}}$, and hence is Noetherian because ${\mathcal{H}_{\operatorname{aff}}}$ is Noetherian (cf. [35, Theorem 5.1] and [23, Theorem 1.4.5]). Since ${\mathcal{H}}_{\chi}$ is finite over ${\mathcal{H}}_{\chi}^{\text{free}}$, it must be Noetherian as well. ∎
We now suppose that $\chi=(J,\xi)$ is a supersingular character of ${\mathcal{H}_{\operatorname{aff}}}$. Given a finite dimensional, irreducible (right) representation $(V,\rho)$ of $\widetilde{\Omega}_{\chi}$ on which $T(\mathbb{F}_{q})$ acts by the character $\xi$, we let $\chi\otimes V$ be the ${\mathcal{H}}_{\chi}$-module whose underlying vector space is $V$ and on which ${\mathcal{H}_{\operatorname{aff}}}$ acts by $\chi$ and $\widetilde{\Omega}_{\chi}$ acts via $\rho$. Then every simple supersingular ${\mathcal{H}}$-module is isomorphic to one of the form $(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$, and furthermore two pairs $(\chi,V)$ and $(\chi^{\prime},V^{\prime})$ give rise to isomorphic simple supersingular ${\mathcal{H}}$-modules if and only if the two pairs are $\widetilde{\Omega}$-conjugate (cf. [35, Theorem 6.18]). Here we say that $(\chi,V)$ and $(\chi^{\prime},V^{\prime})$ are $\widetilde{\Omega}$-conjugate if $\chi^{\prime}=\chi^{\tilde{\omega}}$ for some $\tilde{\omega}\in\widetilde{\Omega}$ and $V^{\prime}\cong V^{\tilde{\omega}}$, where $V^{\tilde{\omega}}$ is the representation of $\widetilde{\Omega}_{\chi^{\tilde{\omega}}}=\tilde{\omega}^{-1}\widetilde{\Omega}_{\chi}\tilde{\omega}$ with underlying vector space $V$ obtained by letting $v\cdot\tilde{\omega}^{-1}\tilde{\omega}^{\prime}\tilde{\omega}:=v\cdot\tilde{\omega}^{\prime}$ for all $v\in V$ and $\tilde{\omega}^{\prime}\in\widetilde{\Omega}_{\chi}$.
@@
We next recall a theorem of Koziol, which describes when simple supersingular modules are trivial in the Gorenstein projective model structure. The statement in loc. cit. doesn’t actually contain (ii) as below, but their proof actually shows the statement as written here.
Theorem 5.4 ([21, Theorem 7.7]).
@@
We now describe ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{\chi}}({\mathfrak{m}})$, where ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ is a simple supersingular ${\mathcal{H}}$-module. Observe first that for any $\mathfrak{n}\in{\operatorname{Mod}}({\mathcal{H}}_{\chi})$ we obtain from the braid relations a direct sum decomposition
$\mathfrak{n}\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}=\bigoplus_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}\mathfrak{n}\otimes T_{\tilde{\omega}}.$
This is in fact a decomposition as an ${\mathcal{H}}_{\chi}$-module. Indeed, for all $h\in{\mathcal{H}_{\operatorname{aff}}}$ and all $x\in\mathfrak{n}$ we have $(x\otimes T_{\tilde{\omega}})\cdot h=x(T_{\tilde{\omega}}hT_{\tilde{\omega}}^{-1})\otimes T_{\tilde{\omega}}$. Furthermore, the group $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\tilde{\omega}}}=\tilde{\omega}^{-1}\widetilde{\Omega}_{\chi}\tilde{\omega}$ naturally acts on $\mathfrak{n}\otimes T_{\tilde{\omega}}$ (cf. 5.3). Note that these actions do not depend, up to isomorphism, on the choice of coset representatives. When $\mathfrak{n}=\chi\otimes V$, the summand $\mathfrak{n}\otimes T_{\tilde{\omega}}$ naturally identifies with $\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}$ as an ${\mathcal{H}}_{\chi}$-module. We therefore get a decomposition
(5.4) ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{\chi}}({\mathfrak{m}})\cong\bigoplus_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}}.$
We finish by recalling that to any character $\xi:T(\mathbb{F}_{q})\to k^{\times}$ we may attach the idempotent
(5.5) $e_{\xi}=\left|T(\mathbb{F}_{q})\right|^{-1}\sum_{t\in T(\mathbb{F}_{q})}\xi(t)T_{t^{-1}}\in k[T(\mathbb{F}_{q})].$
Then we get that $e_{\xi}T_{\hat{s}}=T_{\hat{s}}e_{s\cdot\xi}$ for all $s\in S$, where $s\cdot\xi$ denotes the lift to $W$ of the $W_{0}$-action on characters of $T(\mathbb{F}_{q})$ by conjugation of the argument. If we let $S_{\xi}$ be as in (2.2), then we have $s_{\alpha}\cdot\xi=\xi$ for all $s_{\alpha}\in S_{\xi}$ by [29, Lemma 2.9], and it follows straightforwardly from (2.4) that
(5.6) $(e_{\xi}T_{\hat{s}})^{2}=\begin{cases}-e_{\xi}T_{\hat{s}}&\text{if $s\in S_{\xi}$}\\ 0&\text{otherwise}\end{cases}.$
For later use we record the following immediate consequence of the above:
Lemma 5.5.
Suppose that $\xi:T(\mathbb{F}_{q})\to k^{\times}$ is a character and let $F\subseteq\overline{C}$ be a face such that $S_{F}\subseteq S_{\xi}$. Then $e_{\xi}$ is central in ${\mathcal{H}}_{F}$ and $(e_{\xi}T_{\hat{s}})^{2}=-e_{\xi}T_{\hat{s}}$ for all $s\in S_{F}$.
@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
Our first aim is to understand morphisms between supersingular characters in the homotopy category ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. Fix a face $F\subseteq\overline{C}$. Recall that ${\mathcal{H}}_{F}$ is a Frobenius algebra and in particular selfinjective (cf. [9, Example 4.1(ii)]), and thus the Gorenstein projective model structure on ${\operatorname{Mod}}({\mathcal{H}}_{F})$ has all objects cofibrant and fibrant. In particular, we see from 2.14 that
@@
where two morphisms $f$, $g$ are equivalent under $\sim$ if and only if $f-g$ factors through a projective ${\mathcal{H}}_{F}$-module. We collect the facts we need about the relations between the Gorenstein projective model structures on the categories ${\operatorname{Mod}}({\mathcal{H}}_{F})$, $F\subseteq\overline{C}$, and ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ below:
@@
    $[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}\cong\begin{cases}0&\text{if ${\mathfrak{m}}$ is projective}\\ {\operatorname{Hom}}_{{\mathcal{H}}_{F}}({\mathfrak{m}},\mathfrak{n})&\text{otherwise}.\end{cases}$
@@
    is right Quillen. Moreover, the Gorenstein projective model structure on ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ is the right transfer of the product model structure on $\prod_{F\subseteq\overline{C}}{\operatorname{Mod}}({\mathcal{H}}_{F})$ (i.e. with all classes of morphisms defined termwise) along $\Delta$. In particular, ${\operatorname{pd}}_{{\mathcal{H}_{\operatorname{aff}}}}({\mathfrak{m}})<\infty$ if and only if ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$ is projective for all $F\subseteq\overline{C}$.
@@
For (i), we have that $[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}=0={\operatorname{Hom}}_{{\mathcal{H}}_{F}}({\mathfrak{m}},\mathfrak{n})$ if the simple modules ${\mathfrak{m}}$ and $\mathfrak{n}$ are not isomorphic (cf. (5.7)). If on the other hand ${\mathfrak{m}}\cong\mathfrak{n}$, then by Schur’s lemma we have that ${\operatorname{Hom}}_{{\mathcal{H}}_{F}}({\mathfrak{m}},{\mathfrak{m}})\cong k$ is one-dimensional and hence either its quotient $[{\mathfrak{m}},{\mathfrak{m}}]_{{\mathcal{H}}_{F}}$ is zero or the homotopy relation is trivial. But if $[{\mathfrak{m}},{\mathfrak{m}}]_{{\mathcal{H}}_{F}}=0$ then ${\operatorname{id}}:{\mathfrak{m}}\to{\mathfrak{m}}$ factors through a projective, from which we deduce that ${\mathfrak{m}}$ is projective. We thus get our desired isomorphism.
Part (ii) is immediate from 2.13. For (iii), we obtain as a consequence of (ii) that $\Delta$ preserves fibrations and weak equivalences as those are defined termwise in the target category. Since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}$ is also faithful for each $F\subseteq\overline{C}$ and all objects are fibrant, we in fact get that $\Delta$ reflects fibrations. We are therefore left to show that $\Delta$ has a left adjoint and reflects weak equivalences. For the former, the left adjoint is given by $({\mathfrak{m}}_{F})_{F\subseteq\overline{C}}\mapsto\bigoplus_{F\subseteq\overline{C}}{\mathfrak{m}}_{F}\otimes_{{\mathcal{H}}_{F}}{\mathcal{H}_{\operatorname{aff}}}$. For the latter, by 2.8 it suffices to show that $\Delta$ reflects trivial objects, i.e. that if ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ is such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$ is projective for all $F\subseteq\overline{C}$ then ${\operatorname{pd}}_{{\mathcal{H}_{\operatorname{aff}}}}({\mathfrak{m}})<\infty$. But this is the reverse implication shown in [21, Lemma 4.2]. ∎
5.6(i) has the following immediate consequence which will be helpful later.
@@
    $(\chi^{\prime}(T_{\hat{s}}),\chi^{\prime}(T_{\hat{s^{\prime}}}),\chi^{\prime}(T_{\hat{s^{\prime\prime}}}))$ $=(0,-1,0)$
    and $s^{\prime},s^{\prime\prime}$ are adjacent in the affine Dynkin diagram of $\Phi$. Then ${\operatorname{Ho}}(\Delta)([\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}})$ has dimension 1 over $k$ and doesn’t contain any isomorphisms.
@@
Later we will consider simple supersingular modules for the whole pro-$p$ Iwahori-Hecke algebra ${\mathcal{H}}$, for which the action of $\widetilde{\Omega}$ on supersingular characters comes into play. Suppose that $\omega\in\Omega$ and fix a lift $\hat{\omega}\in\widetilde{\Omega}$. Recalling that the group $\Omega$ normalises $S$, if $\chi$ is any character of ${\mathcal{H}_{\operatorname{aff}}}$ corresponding to a pair $(J,\xi)$ then $\chi^{\hat{\omega}}$ corresponds to the pair $(J^{\hat{\omega}},\xi^{\hat{\omega}})$ where $J^{\hat{\omega}}=\omega^{-1}J\omega$ and $\xi^{\hat{\omega}}(t)=\xi(\hat{\omega}t\hat{\omega}^{-1})$ ($t\in T(\mathbb{F}_{q})$). It follows in particular that $\left|J^{\hat{\omega}}\right|=\left|J\right|$. We therefore see that the two supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ from 5.8(i) are not conjugate under the $\widetilde{\Omega}$-action.
Before proving the Proposition, we first highlight that it immediately implies the following:
Corollary 5.10.
Suppose that $\chi\neq\chi^{\prime}$ are two supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ of infinite projective dimension. Then they are not isomorphic in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. In particular, if $\mathbb{G}$ is semisimple and simply-connected then two supersingular ${\mathcal{H}}$-modules are isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if they are isomorphic in ${\operatorname{Mod}}({\mathcal{H}})$.
Proof.
Any isomorphism $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ would induce an isomorphism ${\operatorname{Ho}}(\Delta(f))$ in $\prod_{F\subseteq\overline{C}}{\operatorname{Ho}}({\mathcal{H}}_{F})$. But it follows immediately from 5.8 that this cannot occur. The last part follows immediately because ${\mathcal{H}}={\mathcal{H}_{\operatorname{aff}}}$ if $\mathbb{G}$ is semisimple and simply-connected. ∎
We start preparing for the proof. Suppose that $\chi$ and $\chi^{\prime}$ are as given in 5.8. Given any $f\in[\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}}$ we obtain a collection $(f_{F})\in\prod_{F\subseteq\overline{C}}{\operatorname{Hom}}_{{\mathcal{H}}_{F}}(\chi,\chi^{\prime})$ by letting each $f_{F}$ be a choice of representative of ${\operatorname{Ho}}(\Delta)(f)_{F}={\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}})(f)$, and this satisfies the property that ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F}}(f_{F^{\prime}})\sim f_{F}$ whenever $F^{\prime}\subseteq\overline{F}$. Note that either $f_{F}\sim 0$ or otherwise the choice of $f_{F}$ must be unique, cf. 5.6(i). Our strategy will be to show that any such collection of morphisms must in fact be zero, up to homotopy, unless we are in the specific circumstances of part (i) of the Proposition. We therefore first need a method for determining whether some $f_{F}$ arising as above is homotopic to zero. To that end, we have the following elementary but key result.
@@
    Assume we have three faces $F_{1},F_{2},F_{3}$ with $F_{i}\subseteq\overline{F_{3}}$ for $i=1,2$ such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective. Then $f_{F_{1}}\sim 0$ if and only if $f_{F_{2}}\sim 0$.
@@
Part (i) is immediate from 5.6(ii) because right Quillen functors preserve right homotopies. For part (ii), assume that $f_{F}\sim 0$ and $f_{F^{\prime}}\not\sim 0$. Then by 5.6(i) we must have ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ and $f_{F^{\prime}}=\lambda{\operatorname{id}}_{\chi}$ for some $\lambda\in k^{\times}$. Then we have ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F}}(\lambda{\operatorname{id}}_{\chi})\sim f_{F}\sim 0$ from which we deduce that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective. For (iii), the condition that $f_{F_{1}}\sim 0$ implies by (i) that $f_{F_{3}}\sim 0$. Since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective it then follows from (ii) that $f_{F_{2}}\sim 0$ as required. We get the converse by swapping $F_{1}$ and $F_{2}$. ∎
In order to apply 5.11, we need some criterion to establish when the restriction of a character of ${\mathcal{H}_{\operatorname{aff}}}$ to some ${\mathcal{H}}_{F}$ is not projective. Note that projectivity always holds when $F=C$ since ${\mathcal{H}}_{C}=k[T(\mathbb{F}_{q})]$ is semisimple. In general, we have the next result below which we gathered from the arguments of Koziol in [21, §7]. To prepare for its proof, we recall that the 00-Hecke algebra over $k$ associated to a finite Coxeter group $W_{\text{gen}}$ with set of Coxeter generators $S_{\text{gen}}$ is the $k$-algebra with generating set $\{H_{s}\mid s\in S_{\text{gen}}\}$ such that the braid relations hold and $H_{s}^{2}=-H_{s}$ for all $s\in S_{\text{gen}}$ (cf. [25, Definition 1.1]). The simple modules for such algebras are all characters and are indexed by the subsets $S_{\text{gen}}$, where for $L\subseteq S_{\text{gen}}$ the associated character $\lambda_{L}$ is defined by $\lambda_{L}(H_{s})=-1$ if $s\in L$ and $\lambda_{L}(H_{s})=0$ otherwise (cf. [25, §3]). When $L=\emptyset$, resp. $L=S_{\text{gen}}$, we call $\lambda_{L}$ the trivial, resp. sign, character.
@@
    If $S_{F}\nsubseteq S_{\xi}$ then ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective.